# Interpretable Responsibility Sharing as a Heuristic for Task and Motion Planning

### v1

@@ 4.3 Optimized Rule Synthesis (ORS)
where $\alpha=0.5$. This optimization process uses validation data to adjust the complexity of the rule set for improved performance. Results from each trial are saved in a structured format allowing for detailed analysis of the rule system’s effectiveness over multiple iterations. The main loop reshuffles and re-evaluates the data multiple times, refining the approach to achieve optimal results based on the performance metrics and saved outcomes. This iterative, data-driven method enhances the adaptability and effectiveness of the rule-based decision-making process by investigating how different rules interact and influence predictions.
In the evaluation phase, these combinations are applied across a dataset, and their outcomes are recorded and sorted by the confidence level of their predictions. If ORS infers the use of auxiliary objects, it also returns predefined usages of these objects, defining sub-problems $\Psi$ for the IRS framework. To elaborate on these predefined usages, if the environment contains a tray that is beneficial to use in a certain scenario, the objects available in the environment first need to be placed on the tray, and then the robot will carry the tray to the final position to complete the task. In this scenario, $\Psi$ is defined in terms of first-order logic as follows: if $s_{0}$ = 4 objects, $s_{g}$ = objects on table, then $\Psi$ = {objects on tray, tray on table} using the carry action of the tray. The complete ORS algorithm is shown in Algorithm 1.
Algorithm 1 Optimized Rule Synthesis
1:Input: Dataset $D$, initial state $s_{0}$, goal state $s_{g}$
2:Output: Set of sub-problems $\Psi$, set of conditions for Responsibility Sharing (RS) $S_{c}$
3:function Optimized Rule Synthesis($D,s_{0},s_{g}$)
4:    $AllRuleConditions\leftarrow$ CARL(D) $+$ RRL(D) $\triangleright$ Rules from CARL and RRL.
5:    $\Psi\leftarrow\emptyset$
6:    for $iteration\leftarrow 1$ to $\beta$ do $\triangleright$ Perform the process $\beta$ times (e.g. $\beta=10$).
7:        $train,val,test\leftarrow$ SplitData($D$)
8:        $GeneratedRules\leftarrow$GeneratingRules($train,AllRuleConditions,iteration$)
9:        Append($OptimalRuleSet$,FindBestMinComponent($val,GeneratedRules$))
10:    end for
11:    $S_{c}\equiv OptimalRuleSet$
12:    if $s_{0}\cup s_{g}\in OptimalRuleSet$ then $\Psi\leftarrow$ ObjectUsage($s_{0}$)
13:    end if
14:    return $\Psi,S_{c}$
15:end function
Briefly, the initial and exhaustive set of rules are obtained from RRL and CARL. The SplitData function randomly splits the data samples into $50\%$ for training, with the remaining $50\%$ split in half between validation and test sets. Starting with a single rule, GeneratingRules finds combinations of these rules, containing $iteration$ many rules in one combination per cycle. This process continues until we find combinations consisting of $\beta$ many rules, which is 10 in this study.
In each iteration, the FindBestMinComponent function is utilized to determine the optimal number of components (or rules) that should be considered when evaluating combinations of rule conditions on the train and validation sets to achieve the best performance in terms of accuracy and interpretability (according to the Balance Score). After finding the optimal set, if the given initial and goal states satisfy the rules generated by the ORS, instructions for using the auxiliary objects defined under sub-problems are found by ObjectUsage. This function formulates the predefined usage of these objects in terms of first-order logic. Finally, ORS returns a set of conditions for Responsibility Sharing $S_{c}$ (equivalent to $OptimalRuleSet$) and a set of sub-problems $\Psi$ for IRS to use.
@@ 4.4 Interpretable Responsibility Sharing as a Heuristic
In this work, we propose a heuristic-based interpretable formulation that guides symbolic task search and leverages the human bias frequently present in sequential decision-making and manipulation tasks. Our heuristics are based on the concept of Responsibility Sharing (RS) and aim to divide the original problem into sub-problems if it enhances the performance of the embodied agent. To implement this, we integrated our approach into Logic-Geometric Programming (LGP)  [4] and used Multi-Bound Tree Search (MBTS)  [65] as the task planning algorithm. This integration leverages the strengths of both LGP and MBTS, ensuring robust and efficient task planning. Importantly, IRS is independent of the underlying search mechanism, making our method orthogonal to existing heuristic-based approaches. IRS can be combined with any search heuristic if it is believed to improve performance on a specific task.
As mentioned in Section 4.3, based on the initial and goal states, $s_{0}$ and $s_{g}$, ORS generates logical rules that dictate the appropriate conditions under which the agent should share responsibilities with auxiliary objects. If the agent decides to use the auxiliary objects, sub-goals are created to satisfy the necessary conditions for their usage. These conditions are object-dependent and naturally occur during their usage. For example, if the agent wants to use a tray to carry objects, the objects need to be placed on the tray first. These preconditions create clear and intuitive sub-goals for the agent to operate effectively.
Motivated by CPG defined in Section 4.2.2, if an agent decides to utilize the auxiliary objects in the environment (i.e., the initial and goal states follow the generated conditions), it will generate a task plan that uses the auxiliary objects by dividing the search into sub-goals defined by the mentioned preconditions. Following the transition function $\delta$, the agent will use MBTS to expand from the initial state $s_{0}$ to ultimately reach the goal state $s_{g}$ while satisfying the sub-goals $s_{i}$, the number of which depends on the number of auxiliary objects. By introducing a sub-goal-based formulation, our approach increases the number of generated plans compared to conventional MBTS if we consider each sequence from a state to a sub-goal as a different plan. However, this naive approach is computationally expensive for a combined task and motion planning approach since the solution time is highly dominated by the geometric calls made by the LGP approach, and these calls should be minimized for efficiency  [16].
Our approach achieves search efficiency by deferring the bounds of Multi-Bound Tree Search (MBTS), as mentioned in Section 3.2. This is necessary given the increased need for nonlinear optimizer calls due to multiple plans generated to satisfy the sub-goals. It prioritizes $P_{\textrm{seq}}$ for optimizing mode-switch sequences (e.g., changing from free space movement to object grasping) and defers $P_{\textrm{path}}$, the complete problem (i.e., considering kinematics, dynamics, and including obstacle and collision checks), until its full realization. While this does not add novelty, it is essential. In the original LGP formulation (Equation 1), these bounds are solved sequentially for a single joint optimization problem, meaning a solution to the lower bounds needs to be found before proceeding to a higher bound. However, in our approach, we introduce auxiliary object-dependent sub-goals formulated as separate plans. Given this abundance, we avoid solving joint motion separately multiple times and instead utilize the found mode-switch sequences for solving a complete motion plan. The adaptation of sub-problems and deferring in MBTS necessitates modifications to the existing LGP formulation.
[…]
@@
Here, we provide the details of our strategy called Mini-LGP, as formulated in Equation 11. The environment offers $m$ distinct auxiliary objects that can share responsibility with the robot. Consequently, the problem can be segmented into a set of sub-problems $\Psi$, where $|\Psi|=2m$. Of these, $m$ are dedicated to fulfilling the necessary conditions for using these objects (e.g., filling the pitcher, moving objects onto a tray), while the remaining $m$ focus on executing shared responsibilities. Each sub-problem is defined and optimized jointly through Mini-LGPs over a problem-specific space. The initial configuration $\bar{x}(0)$ represents the starting state for the first sub-problem or the terminal state from the preceding sub-problem. The set of available actions, $A_{\psi}$, and the set of sub-goal states, $S_{\text{goal}}(g_{\psi})$, are specific to each sub-problem (e.g., the robot cannot fill the glasses using a tray).
Thus, while the core concept of LGP is retained, certain modifications are introduced to ensure compatibility with sub-problem definitions and to integrate human bias through the use of auxiliary objects into the first-order logic definitions. If a solution to Equation 11 exists, our formulation successfully generates a task plan using IRS as a heuristic with relaxed constraints. Now, our formulation will tackle the complete motion problem, $P_{path}$, considering the $P_{path}$ constraints with already computed mode-switch sequences $P_{seq}$. The complete IRS algorithm is shown in Algorithm 2.
@@
6:    Set of sub-problems $\Psi$, set of conditions for RS $S_{c}$leftarrow ORS(D)
@@
[…]
11:               if $P_{seq}$ then $SeqeunceFound\leftarrow true$
12:               end if
13:           until $SequenceFound$
[…]
15:           if $P_{path}$ then $PathFound\leftarrow true$
16:           end if
17:        until $PathFound$
18:    else
19:        repeat
20:           $P_{path}$leftarrow $\textit{solvePath}(s_{0},x_{0},s_{g})$
21:           if $P_{path}$ then $PathFound\leftarrow true$
22:           end if
23:        until $PathFound$
24:    end if
25:    return $P_{path}$
26:end function
In short, ORS is first trained with a counterfactual plan dataset created by CPG. It returns rules that determine the appropriate conditions to employ RS with auxiliary objects, $S_{c}$. Depending on the auxiliary objects, it also returns a set of sub-problems, $\Psi$. These sub-problems are generic to each auxiliary object and require no further attention. We are not concerned with how to use the tray or pitcher, but when to use them. Hence, their utilization is assumed to be already defined based on the environment.
If the conditions generated by ORS are not satisfied, indicating that using auxiliary objects in $s_{0}$ and $s_{g}$ results in a performance decrease, then the LGP solver, solvePath, formulated in Equation 1, is used to find a motion plan. Overall, with its interpretable conditions generated by ORS, IRS guides the TAMP formulation by leveraging the human bias existing in sequential decision-making and manipulation problems.
@@ 5 Experiments
This section primarily validates the effectiveness of IRS as a heuristic for Task and Motion Planning and the interpretability of the underlying decision-making mechanism, ORS, through several aspects. First, we compare our method’s performance to other baselines in completing various household tasks, evaluating both the robot’s effectiveness in minimizing displacement (representing the amount of effort) and the decision-making ability of ORS in determining the optimal time to share responsibility with auxiliary objects. Second, we compare human decision-making with ORS by observing how humans use auxiliary objects while completing a representative set of household tasks that ORS has also been tested on, analyzing the correspondence in object utilization to validate the human bias leveraged by IRS. Third, we present ablation studies on how ORS’s dynamic integration behaves under the interpretability and accuracy trade-off.
@@ 5.1.1 Tasks
Figure 5: Experimental settings used in simulation and human experiments. In the simulation, a north sign is placed to identify the tables mentioned in the task descriptions.
We evaluate IRS as a planning heuristic for task and motion planning in three distinct household tasks: serving, pouring, and handover. All of these tasks are modeled in simulation environments as replicas of real robotic tasks. To infer the usage of auxiliary objects for responsibility sharing with a single model, we utilized an overlapping set of descriptions for the environments, such as $table_{position}$ and $obj_{index}$. Additionally, since one of our main motivations in this paper is leveraging the human bias in sequential decision-making and manipulation problems through the concept of responsibility sharing, we wanted to investigate whether humans also utilize auxiliary objects when they provide a benefit. To do this, we structured a miniature but representative set of experiments based on the serving task and compared human behavior with IRS.
@@
Human Experiments: The development of IRS is notably influenced by intuitive human thinking, making human experiments a crucial aspect of our study to observe human bias. For these experiments, we replicated a portion of the serving task using a varying number of objects (mugs) and a single tray initially positioned on the counter. We asked six human subjects to carry the mugs to their final positions based solely on intuition, without specific guidance on using the available objects. While we are not suggesting that IRS is a model for human sequential decision-making, we believe that observing when humans use these auxiliary objects will strengthen our claim about the existence of human bias in these sequential decision-making problems and how IRS leverages it (Figure 5(d)).
@@ 5.1.2 Baselines
We compare IRS with several baselines to evaluate its effectiveness. First, we compare it with LGP  [4, 68], which does not prioritize using auxiliary objects, and a baseline agent that always prioritizes using auxiliary objects, to show that responsibility sharing depends on the environmental setting and that IRS can improve the existing TAMP formulation. Then, we compare ORS’s decision-making performance with well-established baselines, including Artificial Neural Network (ANN)  [70], XGBoost  [71], Support Vector Machines (SVM)  [72], Logistic Regression  [73], and Decision Tree  [74]. Among these baselines, ANN and Decision Tree represent two opposite ends of the interpretability spectrum, providing a range for comparison. Additionally, the performance of RLL  [25] and CARL  [26] is investigated, enabling comparison with interpretable state-of-the-art methods.
5.1.3 Evaluation Metrics
For baseline comparison and ablation studies, we constructed a dataset using the process explained in Section 4.2.3 and performed randomized split testing, repeated across 5 different random seeds, including all tasks collectively. We then calculated the following evaluation metrics on test scenarios: Effort, Accuracy, Precision, Recall, F1-Score, Confidence, and Interpretability. Effort is measured by the total unit displacements of the robot’s joints in the environment required to complete the task. Accuracy, Precision, Recall, and F1-Score metrics were used to measure decision-making performance in terms of using IRS as a heuristic. Confidence and Interpretability are used to assess the quality and explainability of the proposed model and existing interpretable methods. Findings are reported as the mean and standard deviation across the folds.

### v2

@@ 4.3.4 Iterative Search Procedure
5:for $iteration=1$ to $\beta$ do
6:  $(D_{\text{train}},D_{\text{val}},D_{\text{test}})\leftarrow\text{SplitData}(D)$
7:  $AllRuleConditions\leftarrow\text{CARL}(D_{\text{train}})\cup\text{RRL}(D_{\text{train}})$
8:  $S\leftarrow\emptyset$
9:  repeat
10:   Choose $r\in AllRuleConditions\setminus S$ that maximizes $\mathrm{Score}(S\cup\{r\};D_{\text{val}})$
11:   if $\mathrm{Score}(S\cup\{r\};D_{\text{val}})>\mathrm{Score}(S;D_{\text{val}})$ and $|S|<K_{\max}$ then
12:     $S\leftarrow S\cup\{r\}$ $\triangleright$ Generating Rules
13:   else
14:     break
15:   end if
16:  until no improvement
17:  Append $S$ to OptimalRuleSetsOptimalRuleSets
18:end for
[…]
20:return $S_{c}$
@@ 4.4 Interpretable Responsibility Sharing as a Heuristic
ORS generates logical rules that dictate the appropriate conditions under which the agent should share responsibilities with auxiliary objects. If the agent decides to use the auxiliary objects, sub-goals are created to satisfy the necessary conditions for their usage. These conditions are object-dependent and naturally occur during their usage. For example, if the agent wants to use a tray to carry objects, the objects need to be placed on the tray first. These preconditions create clear and intuitive sub-goals for the agent to operate effectively.
When the ORS conditions are satisfied ($S_{c}$ is true for $s_{0},s_{g}$), the agent adopts the IRS strategy. By introducing a sub-goal-based formulation, our approach increases the number of generated plans compared to conventional MBTS if we consider each sequence from a state to a sub-goal as a different plan. However, this naive approach is computationally expensive for a combined task and motion planning approach since the solution time is highly dominated by the geometric calls made by the LGP approach, and these calls should be minimized for efficiency  [4].
Solving these sub-problems sequentially with full geometric validation is computationally prohibitive due to the high cost of non-linear optimization. To ensure tractability, we utilize the bound-deferral strategy of Multi-Bound Tree Search (MBTS). We prioritize $P_{\textrm{seq}}$ for optimizing mode-switch sequences (e.g., changing from free space movement to object grasping) and defer $P_{\textrm{path}}$, the complete problem (i.e., considering kinematics, dynamics, and including obstacle and collision checks), until its full realization. Given this abundance, we avoid solving joint motion separately multiple times and instead utilize the found mode-switch sequences for solving a complete motion plan. This necessitates modifications to the existing LGP formulation to support segmented optimization, which we term Mini-LGP:
$\hskip 5.69054pt\min_{x,s_{1:K_{\psi}},a_{1:K_{\psi}},K_{\psi}}\hskip 5.69054pt\sum_{\psi=1}^{\Psi}\sum_{k=1}^{K_{\psi}}\int_{T_{k-1}}^{T_{k}}c(\bar{x}(t),s_{k}(t))\,dt$ (9)
@@
The environment offers $m$ distinct auxiliary objects that can share responsibility with the robot. Consequently, the problem can be segmented into a set of sub-problems $\Psi$, where $|\Psi|=2m$. Of these, $m$ are dedicated to fulfilling the necessary conditions for using these objects (e.g., filling the pitcher, moving objects onto a tray), while the remaining $m$ focus on executing shared responsibilities. Each sub-problem is defined and optimized jointly through Mini-LGPs over a problem-specific space. The initial configuration $\bar{x}(0)$ represents the starting state for the first sub-problem or the terminal state from the preceding sub-problem. The set of available actions, $A_{\psi}$, and the set of sub-goal states, $S_{\text{goal}}(g_{\psi})$, are specific to each sub-problem (e.g., the robot cannot fill the glasses using a tray).
If a solution to Equation 9 exists, our formulation successfully generates a task plan using IRS as a heuristic with relaxed constraints. Now, our formulation will tackle the complete motion problem, $P_{path}$, considering the $P_{path}$ constraints with already computed mode-switch sequences $P_{seq}$. The complete IRS algorithm is shown in Algorithm 2.
Depending on the auxiliary objects, it also returns a set of sub-problems, $\Psi$. These sub-problems are generic to each auxiliary object and require no further attention. We are not concerned with how to use the tray or pitcher, but when to use them. Hence, their utilization is assumed to be already defined based on the environment.
If the conditions generated by ORS are not satisfied, indicating that using auxiliary objects in $s_{0}$ and $s_{g}$ results in a performance decrease, then the LGP solver, solvePath, formulated in Equation 1, is used to find a motion plan.
@@
6:  Set of sub-problems $\Psi$, set of conditions for RS $S_{c}$ $\leftarrow$ ORS(D)
@@
10:      $P_{seq}$ $\leftarrow$ $\textit{solveMini-LGP}(s_{0},x_{0},s_{g},\Psi)$
11:      if $P_{seq}$ then
12:       $SequenceFound\leftarrow true$
13:      end if
14:     until SequenceFoundSequenceFound
15:     $P_{path}$ $\leftarrow$ $\textit{solvePathWithSequence}(s_{0},x_{0},s_{g},P_{seq})$
16:     if $P_{path}$ then
17:      $PathFound\leftarrow true$
18:     end if
19:   until PathFoundPathFound
20:  else
21:   repeat
22:     $P_{path}$ $\leftarrow$ $\textit{solvePath}(s_{0},x_{0},s_{g})$
23:     if $P_{path}$ then
24:      $PathFound\leftarrow true$
25:     end if
26:   until PathFoundPathFound
27:  end if
28:  return $P_{path}$
29:end function
4.5 Computational Complexity
In this section, we analyze the computational complexity of the proposed framework, focusing on the tractability of the training phase (ORS) and the inference efficiency of the planning phase (IRS).
We define the following parameters for our analysis: $n$ denotes the number of training instances; $R$ represents the number of candidate atomic rules generated by RRL/CARL; and $L_{\max}$ is the maximum rule length. For the iterative search, $\beta$ is the repetition count and $K_{\max}$ is the maximum size of the final rule set.
4.5.1 Training Complexity
The training process consists of candidate generation followed by rule selection.
1. Candidate Generation (RRL & CARL): RRL training scales linearly with the number of instances and active edges, with a per-epoch cost of $\mathcal{O}(n(b+W))$. CARL complexity is determined by path sampling and attention mechanisms over the knowledge graph, scaling as $\tilde{\mathcal{O}}(S\cdot\ell\cdot H\cdot d_{\text{emb}})$. Both steps are performed once per iteration and remain strictly polynomial.
2. Optimized Rule Synthesis (ORS): To identify the optimal rule set without incurring the exponential cost of exhaustive enumeration ($\mathcal{O}(R^{K_{\max}})$), ORS utilizes a bounded greedy forward selection strategy. In each of the $\beta$ repetitions, the algorithm iteratively adds the single rule $r\in\mathcal{R}$ that maximizes the validation score.
The cost to evaluate one candidate rule on $n$ validation samples is $\mathcal{O}(n\cdot L_{\max})$. Since the algorithm scans at most $R$ candidates for up to $K_{\max}$ steps across $\beta$ repetitions, the total complexity is:
$\mathcal{T}_{\text{ORS}}=\mathcal{O}(\beta\cdot K_{\max}\cdot R\cdot n\cdot L_{\max})$ (10)
This complexity is linear with respect to the candidate pool size $R$ and dataset size $n$. We constrain $L_{\max}\in\{3,4,5\}$, which ensures computational efficiency while enforcing TAMP-level interpretability, where rules remain concise yet descriptive enough to capture task context.
4.5.2 Planning and Inference Efficiency
It is crucial to note that the rule synthesis cost ($\mathcal{T}_{\text{ORS}}$) is a one-time offline expense. During online planning (inference), the agent merely evaluates the pre-generated rule set $S_{c}$.
We compare this online runtime against the Uninformed Search baseline used in our experiments. Uninformed Task and Motion Planning typically explores a state space with branching factor $b_{s}$ to depth $T$, yielding a complexity of $\mathcal{O}(b_{s}^{T})$. By utilizing IRS, the task is decomposed into sub-problems. If IRS identifies a valid sub-goal (e.g., place on tray), the effective search depth for the detailed motion planner is reduced to the sub-goal horizon $T_{sub}\ll T$, or the branching factor is pruned by focusing only on relevant auxiliary objects. Thus, IRS effectively reduces the exponent of the search complexity.
@@ 5 Experiments
This section primarily validates the effectiveness of IRS as a heuristic for Task and Motion Planning and the interpretability of the underlying decision-making mechanism, ORS, through several aspects. First, we compare our method’s performance to other baselines in completing various household tasks, evaluating both the robot’s effectiveness in minimizing displacement (representing the amount of effort) and the decision-making ability of ORS in determining the optimal time to share responsibility with auxiliary objects. Second, we compare human decision-making with ORS by observing how humans use auxiliary objects while completing a representative set of household tasks that ORS has also been tested on, analyzing the correspondence in object utilization to validate the human-centric environmental bias leveraged by IRS. Third, we present ablation studies on how ORS’s dynamic integration behaves under the interpretability and accuracy trade-off.
@@ 5.1.1 Tasks
Figure 5: Experimental Environments and Task Configurations. (a) Serving Task: A mobile PR2 robot transports objects from the counter to various dining tables, deciding whether to use a tray. (b) Pouring Task: The robot distributes water to glasses, choosing between individual filling or using a pitcher. (c) Handover Task: A multi-robot scenario where the mobile agent transfers objects to a stationary manipulator. (d) Human Experiments: A real-world setup replicating the Serving task to validate the correlation between human behavior and IRS decision-making. In simulation images, a North sign provides spatial reference.
We evaluate IRS as a planning heuristic for task and motion planning in three distinct household tasks: serving, pouring, and handover. All of these tasks are modeled in simulation environments as replicas of real robotic tasks. To infer the usage of auxiliary objects for responsibility sharing with a single model, we utilized an overlapping set of descriptions for the environments, such as $table_{position}$ and $obj_{index}$. Additionally, since one of our main motivations in this paper is leveraging the human-centric environmental bias in sequential decision-making and manipulation problems through the concept of responsibility sharing, we wanted to investigate whether humans also utilize auxiliary objects when they provide a benefit. To do this, we structured a miniature but representative set of experiments based on the serving task and compared human behavior with IRS.
@@
Human Experiments: The development of IRS is notably influenced by intuitive human thinking, making human experiments a crucial aspect of our study to observe human-centric environmental bias. For these experiments, we replicated a portion of the serving task using a varying number of objects (mugs) and a single tray initially positioned on the counter. We asked six human subjects to carry the mugs to their final positions based solely on intuition, without specific instructions regarding the use of auxiliary objects. The primary objective of these experiments was to provide empirical grounding for our core hypothesis: that the planning logic modeled in our framework reflects genuine human strategic intuition. To observe this human-centric environmental bias in a condensed and observable form, we designed the experiments to isolate the logical relationship between transport distance and object count from confounding physical variables. We evaluated the alignment by comparing human decisions in these physical trials against the ORS model’s predictions for semantically equivalent scenarios within the simulation (Figure 5(d)).
@@ 5.1.2 Baselines
We compare IRS with several baselines to evaluate its effectiveness. First, we compare it with LGP  [60, 57], which does not prioritize using auxiliary objects, and a baseline agent that always prioritizes using auxiliary objects, to show that responsibility sharing depends on the environmental setting and that IRS can improve the existing TAMP formulation. Then, we compare ORS’s decision-making performance with well-established baselines, including Artificial Neural Network (ANN)  [50], XGBoost  [5], Support Vector Machines (SVM)  [30], Logistic Regression  [9], and Decision Tree  [47]. Among these baselines, ANN and Decision Tree represent two opposite ends of the interpretability spectrum, providing a range for comparison. Additionally, the performance of RRL  [63] and CARL  [29] is investigated, enabling comparison with interpretable state-of-the-art methods.
5.1.3 Evaluation Strategy
For baseline comparison and ablation studies, we constructed a dataset using the process explained in Section 4.2.3 and performed randomized split testing, repeated across 5 different random seeds, including all tasks collectively. In all experiments, we fix the ORS repetition parameter to $\beta=10$, corresponding to the upper end of the stable range identified in our preliminary tuning (we observed that $\beta\in[5,10]$ yields similar validation performance, while smaller values lead to less stable rule sets across runs).
We then calculated the following evaluation metrics on test scenarios: Effort, Accuracy, Precision, Recall, F1-Score, Confidence, and Interpretability. Effort is measured by the total unit displacements of the robot’s joints in the environment required to complete the task. Accuracy, Precision, Recall, and F1-Score metrics were used to measure decision-making performance in terms of using IRS as a heuristic. Confidence and Interpretability are used to assess the quality and explainability of the proposed model and existing interpretable methods. Findings are reported as the mean and standard deviation across the folds.
To specifically assess the reliability of individual symbolic rules generated by the models, we define a Rule Confidence score, $\mathrm{Conf}(r)$. This score estimates the conditional probability that a rule’s prediction is correct given that its antecedent holds. Formally, let $D_{\text{val}}$ denote the validation split, and let $\varphi_{r}(x_{i})\in\{0,1\}$ indicate whether the antecedent of rule $r$ is satisfied on instance $x_{i}$. Let $h_{r}(x_{i})$ be the label predicted by $r$, and $y_{i}$ the ground-truth label. We calculate confidence as the empirical precision on the covered validation instances:
$\mathrm{Conf}(r)=\frac{\sum_{i\in D_{\mathrm{val}}}\mathbb{I}\big[\varphi_{r}(x_{i})=1\wedge h_{r}(x_{i})=y_{i}\big]}{\sum_{i\in D_{\mathrm{val}}}\mathbb{I}\big[\varphi_{r}(x_{i})=1\big]+\varepsilon},$ (11)
where $\varepsilon>0$ is a small constant for numerical stability. It is important to note that $\mathrm{Conf}(r)$ evaluates the rule at the aggregate level (the posterior reliability). It does not assign weights to individual conditions within the antecedent; rather, the relative importance of specific features and conditions is encoded intrinsically during the rule generation phase, via multi-head attention scores in CARL [29] and Gradient Grafting weights in RRL [63]. Therefore, $\mathrm{Conf}(r)$ serves as a post-hoc measure of the resulting rule’s quality rather than a feature importance metric.